GOES 16 Moving to Operational Position 75.2W

GOES-16 Drift and Transition to Operations

January 31, 2018


GOES-16 began drifting to the GOES East operational location of 75.2 degrees west longitude on November 30, 2017. GOES-16 reached 75.2 degrees west on December 11, 2017 and data flow resumed to users on December 14. On December 18, 2017, GOES-16 was declared NOAA’s GOES East operational satellite.

Access GOES-16 operational imagery through the GOES East Image Viewer

On December 14, 2017, GOES-13 direct GVAR was disabled. GOES-13 GVAR and LRIT were relayed through GOES-14 until January 8, 2018, at which time the GOES-13 GVAR relay through GOES-14 GVAR was disabled. GOES-13 ceased transmitting data on January 8 and began drifting to its storage location at 60 degrees west on January 9. The satellite arrived at 60 degrees west on January 31.

Now that the GOES-13 GVAR data stream is permanently deactivated, a number of satellite derived products using GOES-13 data are unavailable. NOAA/NESDIS plans to produce a set of equivalent products from the new GOES East satellite (GOES-16); however, in some cases legacy products will be degraded or discontinued entirely as a result of the loss of GOES East coverage. Therefore, all users should review our product transition status documentation that outlines those changes and/or discontinued products.

Additionally, to facilitate user awareness and effectively prepare for the transition from GOES-13 to GOES-16 GOES East operations, the Office of Satellite and Product Operations (OSPO) has established the an informational webpage on the transition. Users may reference this webpage and refer to ESPC notifications for the latest updates on the transition.

GOES East Transition Plan for Replacement of GOES-13 with GOES-16

Please Note: The following dates are subject to change based on operational needs.

  GOES-13 GOES-16
Date Location Activity Location Activity
Current

75°W (no change)

ESPC Notification of Transition Plan GOES-14 EMWIN and MDL (No SXI or XRS/EUV) are enabled.

89.3°W (no change) ESPC Notification of Transition Plan
11/30/17

Slight drift to eastern edge of GOES East slot at 74.5°W.

All Instruments and Services remain in Operations

Drift start at 1.35°/day moving eastward.

  • Five instruments (ABI, EXIS, GLM, SUVI, SEISS) placed in safe or diagnostic modes with no data capture or distribution.
  • Only MAG continues operating, but with data outages on days of drift start (11/30) and drift stop (12/11).
  • Disable GRB RF, DCS, HRIT/EMWIN, SARSAT.
12/7/17

74.5°W (no change)

Nominal GVAR and LRIT continues and is also rerouted (in duplicate) through GOES-14 GVAR at 105°W. Users should repoint to GOES-14 no later than 12/14/17.

Still drifting

No change

12/11/17

74.5°W (no change)

Beginning 3 weeks of GOES-13 and GOES-16 ‘co-location’ instrument inter-comparisons.

Drift stop at 75.2°W

ABI, EXIS, GLM, SEISS, SUVI commanded to nominal operation modes. ABI INR convergence begins. Still no GRB RF, DCS, HRIT/EMWIN, SARSAT.

12/12/17

74.5°W (no change)

No change

75.2°W (no change)

Perform GOES-16 MAG calibration activity.

12/14/17

74.5°W (no change)

  • Disable DCS, LRIT, SARSAT
  • Disable nominal GVAR and LRIT
  • Rerouted GVAR via GOES-14 still in place until 1/2/18 for purposes of enabling seamless user transition to GOES-16 GRB.

75.2°W (no change)

  • Data distribution resumes for SBN/AWIPS, PDA, GNC-A, CLASS. Activate GRB RF, DCS, HRIT/EMWIN, SARSAT. All data distribution services on for GOES-16.
12/18/17

74.5°W (no change)

No change

75.2°W (no change)

  • GOES-16 products continue validation towards Provisional & Full Validation Maturity levels.
  • GOES-16 products continue validation towards Provisional & Full Validation Maturity levels.
  • Provisional Products (ABI L1b, CMI) no longer require “preliminary, non-operational” data caveat language
1/8/18 74.5°W (no change)
  • Disable GVAR rerouting via GOES-14.
  • Disable GOES-13 EMWIN and MDL.
75.2°W (no change)
1/9/18 Drift start at 0.7°/day moving eastward. 75.2°W (no change)

1/31/18

Drift stop at 60°W

75.2°W (no change)

Details as of January 8, 2018.

Notes: No planned maneuvers / special operations during Thanksgiving and Christmas holidays. No GOES East transition / interruption during Critical Weather Days.

GOES East assignment details which do not appear above:

  1. LZSS data availability and SUVI cal activities are not on this slide because they are not pertinent to the user community.
  2. ANCF/BNCF AWIPS and PDA distribution activities are not on this slide because those notifications are handled separately via NWS and PDA, respectively.
LEGEND:
BLUE Background: indicates GOES East Operational Satellite
Green Background: indicates period of co-located GOES-13 & GOES-16 operations


Contact for further information: Kathryn Mozer at Kathryn.Mozer@noaa.gov
See http://www.ssd.noaa.gov/PS/SATS/messages.html for this and other satellite related messages.